Find the dual basis for the basis {1,1+x,x²-1} of the vector space P3={a0+a1x+a2x²:a0,a1,a2 belongs to R}
1
Expert's answer
2020-03-31T08:39:35-0400

The given vector space is ,

"P_3=\{ a_0+a_1x+a_2x^2:a_0,a_1,a_2\in \R\}" and the given basis is

"\{ 1,1+x,x^2-1\}" .

We known that ,if "B=\{ v_1,v_2,......,v_n\}" is a basis of "V" over "K" and

"\phi_1,\phi_2,......,\phi_n \in V^*" be the linear functionals as defined by


"\phi_i(v_j)=\begin{cases}
1 \ \ if \ i=j \ \\
0 \ \ if \ i\neq j
\end{cases}"


Then "\{ \phi_1,.......\phi_n\}" is a basis of "V^*" called the dual basis of "B" .

According to question,

Let "B=\{v_1=1,v_2=1+x,v_3=x^2-1\}" be a basis of "P_3" .

Now ,any elements of "P_2" is written as linear combination of elements of "B" .

If "a_0+a_1x+a_2x^2 \in P_2" be any arbitrary element then

"a_0+a_1x+a_2x^2=(a_0-a_1+a_2).1+a_1(1+x)+a_2(x^2-1)"

"\implies \ a_0+a_1x+a_2x^2=(a_0-a_1+a_2)v_1+a_1v_2+a_2v_3"

Again, we known that if "U \ and \ V" are two vector space over a field "K" and "\{ v_1,.........,v_n \}" be a basis of "V \ and \ let \ u_1,.......,u_n" be any vector in "U" .Then there exist a unique linear mapping "F:V \rightarrow U \ such \ that \ F(v_1)=u_1,..................,F(v_n)=u_n"

Then ,we can define linear functional as follows,


"\phi_1:P_2\rightarrow \R" defined by "\phi_1(v_i)=\begin{cases}
1 \ if \ i=1 \ \\
0 \ if \ i\neq 1 \\
\end{cases}"


Since,"a_0+a_1x+a_2x^2=(a_0-a_1+a_2)v_1+a_1v_2+a_2v_2"

"\phi_1(a_0+a_1x+a_2x^2)=\phi_1\{(a_0-a_1+a_2)v_1+a_1v_2+a_2v _3\}"

"=(a_0-a_1+a_2)\phi_1(v_1)+a_1\phi_1(v_2)+a_2\phi_1(v_3)"

( Since "\phi_1" is a linear mapping)

i,e, "\phi_1(a_0+a_1x+a_2x^2)=a_0-a_1+a_2"

similarly, "\phi_2:P_2\rightarrow\R" defined by

"\phi_2(v_i)=:\begin{cases}
1 \ if \ i=2 \\
0 \ if \ i\neq 2 \\
\end{cases}"

I,e "\phi_2(a_0+a_1x+a_2x^2)=a_1"

And "\phi_3:P_2:\rightarrow\R" defined by "\phi_3(v_i)=\begin{cases}
1 \ if \ i=3 \\
0 \ if \ i\neq3 \\
\end{cases}"

I,e "\phi_3(a_0+a_1x+a_3x^3)=a_3" .

Therefore "\{\phi_1,\phi_2,\phi_3\}" is the required dual of "Basis \ B" .